# Weakly Supervised Object Segmentation by Background Conditional Divergence

### v1

@@ 4 Results
In the main body, we present results for the real datasets, with appendices containing results on the synthetic Texture+MNIST/FashionMNIST/dSprites C.4 and SAS+dSprites C.3, additional results on real images D, and failure of benchmarks on single-channel images: C.4.1 for synthetic data and D.2 for the SSS dataset. Since our method is weakly supervised, it is noteworthy that our method achieves near-perfect performance (False Positive Rate $<0.5\%$) for background images across all datasets and cluster numbers (except for the SSS dataset, which has its own discussion). As a result, our primary focus will be on the segmentation of foreground objects. When ground truth masks are available we evaluate the segmentation performance on composite images using the following metrics: area-under-the-curve of the receiver operating characteristic (AUCROC), mean intersection of cardinality (IOC) (also known as Jaccard index, which is slightly more conservative than the intersection over union metric) with mask threshold at 0.5, and average precision.
4.1 Side Scan Sonar (SSS) Dataset
We first present visualizations of the input image, ground truth, estimated masking, and thresholded estimated masking for several images in Figure 2. We note that our mask often include more of the shadow, but the ground truth does not account for these shadows. Despite this, our model successfully recovers the ships, providing more precise segmentation in the second and third images. Further examples can be found in Figure 12 where the capturing of shadows is more pronounced. Additionally, we note that because the seabed sand moves, it may cover non-edge parts of the ships, making it reasonable for these regions to be segmented as background. This effect is even more pronounced in Figure 12, which demonstrates two key observations: first, our methodology captures the foreground shadow, an aspect not represented in the ground truth. Second, our model extracts ‘foreground’ objects that are not ships, such as rocks and coarse regions, as illustrated in Figure 14. Finally, in Figure 13 we show more examples of ships and small objects with relatively small shadows.
@@ 4.1 Side Scan Sonar (SSS) Dataset
Figure 2: Segmentation model results on SSS images. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Images sites: Lucinda van Valkenburg.
@@ 4.1.1 Performance against Supervised Baselines
Given that our segmentation method includes additional foreground objects (e.g., rocks), captures shadows, and ignores portions of wrecks covered by sand—elements not considered in the ground truth—we believe a direct comparison with the provided ground truth may not fully reflect the model’s performance. Nonetheless, we benchmark our method against supervised methods selected to reflect the variety of deep learning-based segmentation models used in the vision community: Yang (Yang et al., 2022a), ViT Adapter  (Chen et al., 2022), DeepLabV3 (Chen et al., 2017), HRNet (Wang et al., 2020), U-Net (Ronneberger et al., 2015a), and Salient Object Detection, InSPyReNet (Kim et al., 2023). Following (Sethuraman et al., 2024), we aggregate the images per site and report the IOC for each cite in Table 1.
@@
Table 1: Per-site IoC comparison between our method and supervised benchmark baselines. Bold indicates the best performance per site.
@@ 4.1.2 Unsupervised Segmentation Benchmarks
We attempted to apply unsupervised object segmentation methods introduced by Zou et al. (2023); Chen et al. (2019); Savarese et al. (2021) as benchmarks for the Side Scan Sonar with Shipwreck Objects. We choose ILSGAN (Zou et al., 2023) as it is the best performing method among the generative-based methods, ReDO (Chen et al., 2019) as it makes the minimal assumption about the foreground objects (e.g., it does not assume a minimal object size), and finally the work by Savarese et al. (2021) for its unique attributes (i.e., no pre-trained models, no adversarial training, and no need for data labels).
@@
We use the same pre-processing steps explained in Section 3.4.1 with the same patch size. We adapt the model architectures to accept one channel instead of three. We also use the same hyperparameters used by the authors. We found that the networks produce either trivial masks (i.e., all ones or all zeros) or masks without any meaningful segmentation. The work ReDo by  Chen et al. (2019) converged to either almost all-zero masks or all-ones masks as shown in Figure 16 in the appendix. The foreground mask typically contains either diffuse low-intensity values or small, high-intensity regions. These mask patterns appear to be leveraged by the model to promote statistical dependency—i.e., mutual information—between the input noise vector and the resulting generated image. We use their default setting of generating two masks. Additionally, to ensure a fair comparison, we also tried adding a supervised loss for background-only images, the mean squared loss for the generated mask, which is minimized if the generated mask is all zeros, as in equation 7. Figure 17 shows examples for background and composite patches where our MSE loss drives the foreground mask to be completely all-zeros.
For ILSGAN (Zou et al., 2023), the trained model was able to generate realistic images; however, the corresponding masks almost all zero. In other words, the model degenerates and ignores the possible presence of the shipwreck objects shown in Figure 18 in the appendix. This is not surprising as the ‘foreground’ is very similar to the background.
@@
Finally, we attempted to use the method by Savarese et al. (2021), since it is a model-free method, we applied two schemes: the first one is to input the whole sonar image and the second one use a patch-based scheme with 50% overlap. However, both schemes provide nonsensical masks as expected for an unsupervised method designed to work on color images with the object at the center. Results are shown in Figure 19 in the appendix. Although we show only two examples, this observation is consistent across the testing images.
@@
Admittedly, the hyperparameters selected in prior work, which work when trained on colored natural images, may not be optimal; however, for unsupervised segmentation it is not clear how to conduct the required hyperparameter search. Even with the weak supervision, we hypothesize that these methods may simply be poorly suited for single channel images, such as sonar images, that lack the color and optical cues of foreground and background found in natural imagery. To test this hypothesis, in addition to sonar images, we also benchmark some other works using our synthetic datasets (see C.4).
4.2 CUB Dataset
@@
To qualitatively assess segmentation performance, we show 16 images from the testing set along with their segmentations and the birds masks in Figure 3.
@@
Figure 3: Results of the segmentation model for CUB. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at .5.
@@
In Figure 4, we show a variety of counterfactual images. We notice that the majority of the generated counterfactual images do look like real images. However, there are interesting cases where the masking network, in addition to birds, captures other objects. For example, it captures tree branches (e.g., third row, 6th and 7th columns), rocks (e.g., third row, 14th column), wires (first row, last column), cars (first row, 8th column). These objects can be still considered as foreground, so capturing them is not surprising. Also, we can see some unexplained failure such as the 10th and 12th pictures in the first row, where for the former the counterfactual does not appear to differ from the input image and for the latter, the counterfactual does not look like anything in the CUB dataset.
@@
To delve into more examples, we show the generated counterfactual images using the corresponding source composite images (first column) and the source background images (first row) in Figure 20 in the appendix.
For quantitative results, we show in Table 2 the IOC metric for different cluster numbers under AE+K-means and CL+K-means settings (other metrics are in Table 6).
Table 2: Masking Network Performance (IoC) Comparison of AE+K-means and CL+K-Means based Background Clustering.
@@ 4.2.1 Benchmark Comparison
We also include a comparison between our method and other related work that does not require pretrained networks as in Table 3. Here, we use a 5 cluster method. The model is different than the one previously discussed, as the benchmark (Chen et al., 2019) uses more training data, i.e., 10,000 instances compared to the approx. 5,000 instances used in Table2. With the additional training data, our IoC performance increases by 6 percentage points.
Table 3: IoC scores on the CUB dataset across different methods. Bold indicates the best score. From left to right, each work corresponds to Singh et al. (2019), Chen et al. (2019),Bielski & Favaro (2019),Hung et al. (2019),Savarese et al. (2021),Yang et al. (2022b), He et al. (2022), Zou et al. (2023), respectively.
@@
ReDO PerturbGAN SCOPS Savarse et al. Yang et al. GANSeg ILSGAN Ours
@@ 5 Discussion
We introduced a framework for weakly-supervised foreground object segmentation within a dataset comprising both foreground and background images. Our approach leverages background images to formulate a straightforward loss function that avoids degenerate solutions and offers a more straightforward training scheme compared to generative and ViT-based models. Our approach can be applied without relying on large models pretrained on natural imagery. Methods based on such models often cannot be directly applied to other data modalities. However, our approach can be directly applied to other modalities such as sonar object data. The distinctive combination of our dataset and framework sets our work apart.
We demonstrate the limitations of related work by applying two works by Bielski & Favaro (2019) and Xie et al. (2022) to our synthetic datasets. Both methods failed to converge to a solution where the segmentation results are acceptable. This is due to the lack of a well-pretrained model (for the work done by Xie et al. (2022)) and having diverse classes of foreground objects (for the work done by  Bielski & Favaro (2019)).
In addition to sonar data, we also demonstrate relatively good performance on natural images using the CUB dataset. During training of the masking network, random patches are taken from the input image to assign a background cluster for the conditional divergence loss. This demonstrates that our approach can also be used for datasets composed only of composite images (as long as the images are not tightly cropped around the objects).
@@
The main limitation of our approach is the reliance on an appropriate background clustering. Although we observe that synthetic and CUB dataset achieves relatively similar performance with a different number of clusters constructed under varying settings, we believe the quality of the background clusters is crucial for our method to work. This is more pronounced in sonar dataset, where we have noticed a failure mode for K-means in which it assigns the vast majority of instances to a single cluster. In this case, we had to discourage it from doing so by using Sinkhorn algorithm. Selecting the correct approach to clustering can be a challenging task. Notably, the clustering does not have to correspond to true background classes.

### v2

@@ 6 Results
We firstly present results for the real datasets: AI4Shipwrecks in Section 6.1, and SAS-Clutter in Section 6.2, and CUB-200-2011 in Section 6.3. Then we present the results of the synthetic objects on real sonar images SAS+dSprites in Section 6.4, and results on synthetic composite images Textures+MNIST/FashionMNIST/dSprites in Section 6.5. In the appendix, we show additional results on real images Appendix D, failure of benchmarks on single-channel images: Appendix C.4 for synthetic data and Appendix D.1.1 for the AI4Shipwrecks dataset.
Since our method is weakly supervised, it is noteworthy that our method achieves near-perfect performance (False Positive Rate $0.5\%$) for background images across all datasets and cluster numbers (except for the AI4Shipwrecks dataset, which has its own discussion). As a result, our primary focus will be on the segmentation of foreground objects. When ground truth masks are available we evaluate the segmentation performance on composite images using the following metrics: area-under-the-curve of the receiver operating characteristic (AUCROC), mean intersection of union (IoU) (also known as Jaccard index) with mask threshold at 0.5, and average precision.
6.1 AI4Shipwrecks
We first present visualizations of the input image, ground truth, estimated masking, and thresholded estimated masking for several images in Figure 2. We note that our masks often include more of the shadow, but the ground truth does not account for these shadows. Despite this, our model successfully recovers the ships, providing more precise segmentation in the second and third images. Further examples can be found in Figure 14 where the capturing of shadows is more pronounced. Additionally, we note that because the seabed sand moves, it may cover non-edge parts of the ships, making it reasonable for these regions to be segmented as background. This effect is even more pronounced in Figure 14, which demonstrates two key observations: first, our methodology captures the foreground shadow, an aspect not represented in the ground truth. Second, our model extracts ‘foreground’ objects that are not ships, such as rocks and coarse regions, as illustrated in Figure 15. Finally, in Figure 16 we show more examples of ships and small objects with relatively small shadows.
@@ 6.1 AI4Shipwrecks
Figure 2: Segmentation model results on AI4Shipwrecks. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Images sites: Lucinda van Valkenburg.
@@ 6.1.1 Performance against Supervised Baselines
Given that our segmentation method includes additional foreground objects (e.g., rocks), captures shadows, and ignores portions of shipwrecks covered by sand—elements not considered in the ground truth—we believe a direct comparison with the provided ground truth may not fully reflect the model’s performance. Nonetheless, we benchmark our method against supervised methods selected to reflect the variety of deep learning-based segmentation models used in the vision community: Yang (Yang et al., 2022a), ViT Adapter  (Chen et al., 2022), DeepLabV3 (Chen et al., 2017), HRNet (Wang et al., 2020), U-Net (Ronneberger et al., 2015a), and Salient Object Detection, InSPyReNet (Kim et al., 2023). Following Sethuraman et al. (2024), we aggregate the images per site and report the IoU for each cite in Table 2.
@@
Table 2: Per-site IoU comparison between our method and supervised benchmark baselines on the AI4Shipwrecks dataset. Bold indicates the best performance per site.
@@ 6.1.2 Unsupervised Segmentation Benchmarks
We attempted to apply unsupervised object segmentation methods as benchmarks. We choose ILSGAN (Zou et al., 2023) as it is the best performing method among the generative-based methods, ReDO (Chen et al., 2019) as it makes the minimal assumption about the foreground objects (e.g., it does not assume a minimal object size), and finally the work by Savarese et al. (2021) for its unique attributes (i.e., no pretrained models, no adversarial training, and no need for data labels). For all, we use the same pre-processing steps explained in Section 4.1 with the same patch size.
@@
For ILSGAN and ReDO, we adapt the model architectures to accept one channel instead of three, and use the same hyperparameters as in the original works. We found that the networks produce either trivial masks (i.e., all ones or all zeros) or masks without any meaningful segmentation. The work ReDo by Chen et al. (2019) converged to either almost all-zero masks or all-ones masks as shown in Figure 18 in the appendix. The foreground mask typically contains either diffuse low-intensity values or small, high-intensity regions. These mask patterns appear to be leveraged by the model to promote statistical dependency—i.e., mutual information—between the input noise vector and the resulting generated image. We use their default setting of generating two masks. Additionally, to ensure a fair comparison, we also tried adding a supervised loss for background-only images, the mean squared loss for the generated mask, which is minimized if the generated mask is all zeros, as in equation 7. Figure 19 shows examples for background and composite patches where our MSE loss drives the foreground mask to be completely all-zeros. For ILSGAN (Zou et al., 2023), the trained model was able to generate realistic images; however, the corresponding masks almost all zero. In other words, the model degenerates and ignores the possible presence of the shipwreck objects shown in Figure 20 in the appendix. This is not surprising as the ‘foreground’ is very similar to the background.
@@
Finally, we attempted to use the method by Savarese et al. (2021); since it is a model-free method, we applied two schemes: the first inputs the whole sonar image and the second uses a patch-based scheme with 50% overlap. However, both schemes provide nonsensical masks as expected for an unsupervised method designed to work on color images with the object at the center. Results are shown in Figure 21 in the appendix. Although we show only two examples, this observation is consistent across the testing images.
@@
Admittedly, the hyperparameters selected in prior work, which work when trained on colored natural images, may not be optimal; however, for unsupervised segmentation it is not clear how to conduct the required hyperparameter search. Even with the weak supervision, we hypothesize that these methods may simply be poorly suited for single channel images, such as sonar images, that lack the color and optical cues of foreground and background found in natural imagery. To test this hypothesis, in addition to sonar images, we also benchmark other unsupervised segmentation methods using our synthetic datasets (see Appendix C.4).
6.2 SAS-Clutter
Figure 3 visualizes a random sampling of the test set seafloor clutter objects, their learned masks, and the resulting pixel-wise multiplication of the mask and the object images. In each case, the model successfully masks the bright object highlight. Furthermore, in cases where a shadow exists, it can be seen that the masks extend to encompass them as well. However, while the highlight masks tend to be dilated in size, encompassing the entire object, the shadow masks tend to be slightly eroded. As the vast majority of the object images have no ground truth segmentation mask, comparison with supervised methods is impossible. As with the AI4Shipwrecks dataset, we attempted to compare performance with the three unsupervised approaches ILSGAN (Zou et al., 2023), ReDO (Chen et al., 2019), and the in-painting method by Savarese et al. (2021). Both GAN approaches (ILSGAN and ReDO) failed to produce usable masks. The in-painting approach by Savarese et al. was able to produce reasonable masks; however, at convergence, they covered only the brightest highlights of each image. We computed the IoU over approximately $200$ hand-drawn ground truth masks for our approach and the in-painting approach by Savarese et al.. For the in-painting approach we report both the best IoU over all iterations and the IoU at convergence. Our approach produced an IoU $0.46\pm 0.18$ while the in-painting approach gave $0.36\pm 0.14$ and $0.21\pm 0.15$.
Figure 3: Segmentation model results on SAS-Clutter object images. Each column consists of a object snippet image, its estimated segmentation mask, and the masked object.
6.3 CUB-200-2011
@@
To qualitatively assess segmentation performance on CUB, we show 16 images from the testing set along with their segmentations and the birds masks in Figure 4.
@@
Figure 4: Results of the segmentation model for CUB. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at 0.5.
@@
In Figure 5, we show a variety of counterfactual images. We notice that the majority of the generated counterfactual images do look like real images. However, there are interesting cases where the masking network, in addition to birds, captures other objects. For example, it captures tree branches (e.g., third row, 6th and 7th columns), rocks (e.g., third row, 14th column), wires (first row, last column), cars (first row, 8th column). These objects can be still considered as foreground. Also, we can see some unexplained failure such as the 10th and 12th pictures in the first row, where for the former the counterfactual does not appear to differ from the input image and for the latter, the counterfactual does not look like anything in the CUB dataset.
@@
To delve into more examples, we show the generated counterfactual images using the corresponding source composite images (first column) and the source background images (first row) in Figure 6.
For quantitative results, we show in Table 3 the IoU metric for different cluster numbers under AE+K-means and CL+K-means settings (other metrics are in Table 7).
Table 3: Masking Network Performance (IoU) for Different Background Clusterings.
@@ 6.3.1 Benchmark Comparison
We also include a comparison between our method and other related work that does not require pretrained networks. Here, we use 5 clusters. The model is different than the one previously discussed, as the benchmark (Chen et al., 2019) uses more training data, i.e., 10,000 instances compared to the approximately 5,000 instances used in Table 3. As shown in Table 4, with the additional training data, our IoU performance increases by 6 percentage points; however, our approach is outperformed by most of the models.
Table 4: IoU scores (%) on the CUB dataset across different methods. Bold indicates the best score. From left to right, each work corresponds to Singh et al. (2019), Chen et al. (2019),Bielski & Favaro (2019),Hung et al. (2019),Savarese et al. (2021),Yang et al. (2022b), He et al. (2022), Zou et al. (2023), respectively.
@@
ReDO
PerturbGAN
SCOPS
Savarse et al.
Yang et al.
GANSeg
ILSGAN
Ours
@@
Our model’s modest performance is characteristic of its objective: generating plausible counterfactuals. The masking network finds a solution not aligned to the ground truth due to three main reasons: it can create plausible counterfactuals by taking portions of images that create bird-like silhouettes, as seen in Figure 6 rows 4, 6, 9, 10, and 11; it can ignore portions of birds bodies that can vary in color, such that when blended with background they are still reasonable representations (perhaps of different species); and thirdly, the strong contextual dependency between foreground objects and their backgrounds. When a strong dependency exists, a cleanly segmented foreground object often creates a statistically improbable image when placed in a new context, resulting in a high loss. Instead of this, our model strategically finds masks that include ‘contextual anchors’ from the source background to preserve local realism and create a more plausible final image. For instance, consider a task where the source image is a bird native to a dense forest, and the target distribution is characterized by backgrounds that are mostly open sky with a few branches. To minimize the statistical distance, the model will not just segment the bird; it will segment the ‘bird-on-branch’ unit together. A forest bird floating against a plain sky is a statistical anomaly, but by including the branch, the model preserves a high-probability visual pattern. Similarly, for the same target distribution, creating a silhouette of a flying bird out of any composite image will create a plausible counterfactual with the open sky background. Together, these masks create a counterfactual distribution that better align with the target distribution, thereby achieving a lower loss, but do not align with the ground truth mask of the bird alone. While this masking behavior leads to imperfect segmentation on CUB, it highlights our method’s key strength—it is ideally suited for domains where the foreground and background are decoupled. As shown for the sonar datasets, man-made objects like shipwrecks (artificial reefs, abandoned crab/lobster traps, or unexploded ordinances or mines) do not correlate with the surrounding seabed. In this setup, where contextual priors are not a factor, our method excels, proving competitive even with supervised approaches.
6.4 SAS+dSprites
We present both quantitative and qualitative results for sonar dataset that combines real-world seafloor SAS imagery from Cobb & Zare (2014) with synthetic objects from the dSprites dataset. The seafloor contains six texture categories: sand ripple, rock, coral, sand, mud, and seagrass. Figure 7 illustrates the segmentation performance, demonstrating that the proposed method generates high-quality segmentation masks. For this dataset, the method does not produce ‘sharp’ masks. Table 5 provides the performance metric values for different numbers of background clusters. The results indicate that the best performance is achieved when using four clusters.
Figure 7: Results of the segmentation model for synthetic objects on real sonar images. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at 0.5.
Table 5: Performance across Different Numbers of Clusters for Synthetic Objects on Real Sonar Images
Clusters Avg. Precision IoU AUC-ROC
3 60% 30% 90%
4 82% 57% 97%
5 22% 20% 87%
6.5 Textures+MNIST/FashionMNIST/dSprites
We provide both quantitative and qualitative results for three datasets: MNIST, Fashion MNIST, and dSprites. Initially, we trained the segmentation model using true clusters, defined by the texture label. Subsequently, we trained the segmentation model based on AE clustering. We begin by presenting quantitative results, followed by showcasing qualitative outcomes for the best-performing models on each dataset. We show eight testing images, the estimated masks, the true masks, and the thresholded estimated mask based on $0.5$ in Figure 8 using 10 clusters obtained from the AE+K-Means.
Figure 8: Results of the segmentation model for Fashion MNIST (left), MNIST (middle), and dSprites (right). Top to bottom: input images, estimated masks, true masks, and estimated masks thresholded at 0.5
Figure 9 shows the quantitative performance results for the three datasets with respect to the number of clusters from K-means along with performance using the true texture classes indicated by stars. Qualitative results for the masks using the true texture classes in provided in Figure 11 in the appendix.
Figure 9: The percentage of Average Precision, AUCROC, IoU values (y-axis) for the four datasets, as indicated by the legend. The x-axis represents the chosen number of clusters for the K-means model based on AE/CLE features. The star symbols above the x-axis values, specifically at the position of 10, emphasize the Average Precision and AUCROC corresponding to the ground truth clusters.
The performance of the model on Fashion MNIST, which has relatively large objects with polygonal-like shapes, is high with all clusters performing relatively the same. The model performs similarly on MNIST when utilizing any of the estimated clusters, except for a slight drop when using the 10 clusters. We notice, however, that the performance of the true clusters is worse when compared with the estimated ones, which is evident in the qualitative results as well (see Figure 11 in the appendix). The model’s performance on the dSprites dataset is comparatively lower, which we primarily attribute to the small size of objects and intensity variation. Despite this, the qualitative results depicted in Figure 8 show that as in the case of sonar data, the masks are not ‘sharp’. It is noteworthy that the mask size is proportional to the object size. Notably, when employing AE-based clustering with a cluster number equal to 10, the dataset achieves the highest average precision at $83\%$, compared to $60\%$ when using true clusters.
To further discuss the reason for the superior performance for estimated clusters versus the true texture classes, we show the clustered background using AE+K-means along with the corresponding true clusters in Figure 10 in the appendix. We notice that the K-means clustering creates clusters out of the same true cluster, and each cluster has a different phase. For example, in the case of 5 clusters, the brick cluster is divided into two out-of-phase clusters. It also groups together distinct true backgrounds into one cluster. Another example, in the case of 5 clusters, the first row, which represents a cluster, shows a mix of two true clusters. Consequently, images with essentially the same content but differing phase are not assigned to the same cluster.
@@ 7 Discussion
We introduced a framework for weakly-supervised foreground object segmentation within a dataset comprising both foreground and background images. Our approach leverages background images to formulate a straightforward loss function that avoids degenerate solutions and offers a more straightforward training scheme compared to generative and ViT-based models. Our approach can be applied without relying on large models pretrained on natural imagery. Methods based on such models often cannot be directly applied to other data modalities. In contrast, our approach can be directly applied to other modalities such as sonar data, where the image encodes intensity of the echo. The distinctive combination of our dataset and framework sets our work apart.
We demonstrate the limitations of related work by applying two works by Bielski & Favaro (2019) and Xie et al. (2022) to multiple synthetic datasets. Both methods failed to converge to a solution where the segmentation results are acceptable. This is due to the lack of a well-pretrained model (for the work done by Xie et al. (2022)) and having diverse classes of foreground objects (for the work done by  Bielski & Favaro (2019)).
In addition to sonar data, we also demonstrate relatively good performance on natural images using the CUB dataset, but note that counterfactual generation in cases with strong object-background dependence limits performance. During training of the masking network, random patches can be taken from the input image to assign a background cluster for the conditional divergence loss. This demonstrates that our approach can also be used for datasets composed only of composite images (as long as the images are not tightly cropped around the objects).
We have assumed only access to composite images and background-only images. If we also have access to the ground truth masks, one can incorporate ‘cut, paste and learn’ approaches that use known ground truth masks to help train segmentation models (Dvornik et al., 2018; Dwibedi et al., 2017; Fang et al., 2019) as additional synthetic supervision to be combined with weakly supervised real cases. However, we note that the location selection problem involved in ‘cut, paste and learn’ may not be as crucial in image domains such as sonar. Nonetheless, the various blending techniques by Dvornik et al. (2018) could alleviate some of the artifacts in counterfactuals.
@@
The main limitation of our approach is the reliance on an appropriate background clustering. We observed that each dataset achieves optimal performance with a different number of clusters, constructed under varying settings. Selecting the correct number of clusters is a challenging task, but could be guided by visual inspection of images assigned to potential clusters. Notably, the optimal clustering does not have to correspond to true background classes. On the dSprites synthetic dataset, using the clusters actually outperformed the background classes corresponding to the texture class.
8 Conclusion
We have investigated a novel technique for weakly supervised segmentation based on the assumptions that an object’s appearance is independent of the background imagery and there exists sufficient variety across clusters of backgrounds. Given these assumptions, we can avoid a trivial solution when seeking a masking that creates counterfactuals that superimpose a masked object obtained on a different background by minimizing a divergence conditioned on the background cluster. The method, which does not use any pretrained network, performs well on real-world sonar images of shipwrecks and synthetic images where the assumptions hold and other unsupervised and weakly supervised segmentation methods fail. It achieves modest performance on natural images where only the second assumption holds. Future work can investigate possibilities for semi-supervised approaches that incorporate limited ground-truth masks.